# Nonparametric Estimation of Matching Efficiency and Mismatch in Labor Markets via Public Employment Security Offices in Japan, 1972-2024

### v2

@@ 1.1 Related Literature
Second, this paper contributes to the literature on mismatch (Şahin et al. 2014). The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al. (2016). In the context of Japanese labor markets, several papers estimate mismatch at geographical- and occupation-level using Cobb-Douglas specification (Shibata (2013, 2020), Kawata et al. (2016), Higashi (2018), Kawata (2019), Higashi (2020, 2021), Higashi and Sasaki (2023)). This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index. Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012) and Dubé et al. (2012) to calculate the geographical- and occupation-level mismatch index.
@@ 3.2 Mismatch
4 Monte Carlo simulation
  • How large is the bias of matching efficiency from the Cobb Douglass specification?
  • How large is the bias of mismatch from the Cobb Douglass specification?
  • Mismatch literature:
    • mismatch Shimer (2007)
    • mismatch and large firms (Eeckhout 2018)
    • Classical Jackman and Roper (1987)
@@ 7 Conclusion
Acknowledgments
I thank Fuhito Kojima, Kosuke Uetake, Akira Matsushita, Kazuhiro Teramoto, Ryo Kambayashi, Higashi Yudai for their valuable advice. This work was supported by JST ERATO Grant Number JPMJER2301, Japan.

### v3

@@ 1.1 Related Literature
Second, this paper contributes to the literature on labor market mismatch, building on the foundational work of Şahin et al. (2014) extending the earlier analysis by Jackman and Roper (1987). In the static matching model with multiple sectors, Jackman and Roper (1987) demonstrate that aggregate hires are maximized when unemployment is distributed across sectors in a way that equalizes sectoral labor-market tightness. Şahin et al. (2014) extend this framework by incorporating a dynamic, stochastic environment with multiple sources of heterogeneity, providing a methodology to construct counterfactual measures of unemployment in the absence of mismatch. Their focus is on mismatch unemployment, defined as unemployed workers searching in the “wrong” sector. [^2] Note that a distinct body of literature, such as Eeckhout and Kircher (2011), uses the term “mismatch” to describe employed individuals in suboptimal jobs, where the alignment between worker skills and firm capital is not ideal. Eeckhout (2018) provides an overview of this literature, identifying three primary causes of mismatch: search frictions (Eeckhout and Kircher 2011), stochastic sorting (Chade and Eeckhout 2016), and multidimensional types (Lindenlaub 2017). In addition, Guvenen et al. (2020) explores skill mismatch, which quantifies the gap between the skills required by an occupation and the abilities a worker possesses. Another approach, as in Shimer (2012), takes a flow perspective by decomposing unemployment volatility into transition rates between four labor force states: (i) permanent-contract employment, (ii) temporary-contract employment, (iii) unemployment, and (iv) non-participation, providing a foundation for calculating mismatch measures based on employment status.
The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al. (2016). In the context of Japanese labor markets, several papers estimate mismatch at geographical- and occupation-level using Cobb-Douglas specification (Shibata (2013, 2020), Kawata et al. (2016), Higashi (2018), Kawata (2019), Higashi (2020, 2021), Higashi and Sasaki (2023)). This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index. Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012) and Dubé et al. (2012) to calculate the geographical- and occupation-level mismatch index.